Determine the equivalent damping coefficient for the following nonlinear damper:

F) = Ci + C3.13

where C1 = 5 N s/m and C3 = 0.6 N s3 /m3 .

Note: the damper is to be operated around a speed of 5 m/s.
Answer #1

Solution:

The relation between force and velocity is given by

F) = Ci + C3.13

The equivalent damping coefficient is given by,

dF d C = dide = . (C1i + C3.2%) = C1 +3C3.

Substitute,

C = 5 Ns/m, C3 = 0.6 Ns/m, i = 5 m/s

Therefore,

C=C1 +33.22 = 5+3(0.6) (5) = 14 Ns/m